OpenStudy (anonymous):

Melissa and Kyle went on a trip to visit relatives. On the way there, Melissa drove and averaged about 34 mph. On the way home, Kyle drove and averaged about 50 mph. If the total drive time was 21 hours, how long did Melissa drive? How long did Kyle drive?

OpenStudy (anonymous):

Melissa drove 714 miles and Kyle drove 1050 miles

OpenStudy (kropot72):

Let the distance driven by each person be d miles. Melissa's time = d/34 hours ...............(1) Kyle's time = d/50 hours ....................(2) \[\frac{d}{34}+\frac{d}{50}=21\ ..........(3)\] Solving equation (3) gives d = 425 miles. If you plug this value for d into equations (1) and (2) you will find how long each person drove.